# Chiara Amorino, Vytautė Pilipauskaitė

### v2

@@ 7 Proof technical results
having also used the Lipschitz continuity of $a$ and Lemma 5.1(ii), (iii). A similar bound holds true for the moments of $\alpha_{j+1}^{i}$ which, together with (7.2) and (27), provides
@@
By (7.2) we conclude that $\mathbb{E}[|\mathbb{E}_{j\Delta}[\tilde{\beta}^{i}_{j}\tilde{\xi}^{i}_{j}]|^{k}]\leq C\Delta^{\frac{3}{2}k}$, where $C$ does not depend on $i,j,n,N$. The same argument works for the $k$-th absolute moment of $\mathbb{E}_{j\Delta}[\beta^{i}_{j+1}U^{i}_{j}]$. It remains to deal with
@@
In order to prove (33) for every $\theta$ we use [30, Lemma 9]. We note that $F_{j,k}(\theta)$ is $\mathcal{F}^{N}_{(j+2)\Delta}$-measurable, since so is $\tilde{X}_{(j+1)\Delta}^{i}$. Therefore, we apply [30, Lemma 9] separately to the sum of $F_{j,k}(\theta)$ over even $j$ and that one over odd $j$. However, if we gather all the terms back, then it suffices to prove that, for $k=1,2$,
@@
In order to prove the convergence on the LHS of (43) for every $\theta$, we use [30, Lemma 9] in the same way as in the proof of (33): it suffices to prove that
@@
$\sum_{j=2}^{n-2}F_{j,1}\xrightarrow{d}N(0,\bar{\Pi}((fa)^{2})),\qquad\sum_{j=1}^{n-2}F_{j,2}\xrightarrow{\mathbb{P}}0,\qquad F_{0}\xrightarrow{\mathbb{P}}0.$ (46)
@@
The second relation in (46) follows by [30, Lemma 9] if we prove that both
@@
$R^{N}_{n,1}\xrightarrow{d}N(0,\bar{\Pi}(f^{2}a^{4})),\qquad\sum_{k=2}^{5}R^{N}_{n,k}\xrightarrow{\mathbb{P}}0.$ (51)
We first study the convergence of $R^{N}_{n,1}$. We recall that $U^{i}_{j}=\tilde{\xi}^{i}_{j}+\xi^{i}_{j+1}$, where $\tilde{\xi}^{i}_{j}$, $\xi^{i}_{j+1}$ are $\mathcal{F}^{N}_{(j+1)\Delta}$-, $\mathcal{F}^{N}_{(j+2)\Delta}$-measurable, respectively, and independent, and $N(0,\frac{1}{3})$-distributed. We rewrite
@@
We get that $\mathbb{E}[|\rho_{0}|]\leq C(N\Delta)^{\frac{1}{2}}$, which tends to 0, while the sum of $\rho_{j}$ over $j=1,\dots,n-2$ is the main term. In order to prove that this sum converges in distribution to $N(0,\bar{\Pi}(f^{2}a^{4}))$, we use a central limit theorem for martingale difference arrays (see, e.g. Theorems 3.2 and 3.4 of [42]). In particular, we need to prove that
@@
follows by [30, Lemma 9] (see also (44)) and by Proposition 5.8. This concludes the proof of (53). Finally, (54) follows from the bound $C\Delta\to 0$ on the $L^{1}$-norm of its LHS (see also (49)).
@@
Furthermore, application of [30, Lemma 9] implies that
@@
by $C\Delta$ and $C\Delta$, respectively. Similarly, application of [30, Lemma 9] implies that $R^{N}_{n,3}\xrightarrow{\mathbb{P}}0$ as $N\Delta\to 0$. Also, it is easy to check that $\mathbb{E}[|R^{N}_{n,4}|]\leq C(N\Delta)^{\frac{1}{2}}$ tends to 0. At last, let us prove that $R^{N}_{n,5}\xrightarrow{\mathbb{P}}0$ as $N\Delta\to 0$. By applying the $L^{1}$ argument again, the proof reduces to that of
@@
as $N\Delta\to 0$, which in turn follows by [30, Lemma 9], because using Proposition 5.4(i) we can bound the $L^{1}$-norm of
@@
In the proof below, we will follow the formalism introduced in Section 2.3.2 of [68]. Note that since $A_{2k}=(0,\dots,a^{(k)}(Z_{t}),\dots,0)$ and $a^{(k)}(Z_{t})$ is bounded away from zero, the vectors $(A_{2k})_{k=1,\dots,N}$ are independent and span $\mathbb{R}^{N}$. Our goal is to show that, in order to generate $\mathbb{R}^{2N}$, it is necessary to also consider the vectors $([A_{0},A_{2k}])_{k=1,\dots,N}$.
In our framework, the sparsity of the matrix $A$ significantly simplifies the analysis. Recall that
$A_{0}=B-\frac{1}{2}\sum_{l=1}^{2N}A_{l}^{\nabla}A_{l},$
and observe that this reduces to
@@
since $A_{l}^{\nabla}A_{l}=0$ for odd $l$. Thus, only the even-indexed terms contribute, which simplifies the structure of $A_{0}$. Moreover, according to the definition of covariant derivative as in (11),
$A_{2l}^{\nabla}A_{2l}=\sum_{i,j=1}^{2N}A^{i}_{2l}\partial_{i}A_{2l}^{j}\partial_{z_{j}},$
where $\partial_{z_{j}}=1_{j=2k-1}\partial_{y^{k}}+1_{j=2k}\partial_{x^{k}}$. Recall also that $\forall i\neq 2l$ it is $A^{i}_{2l}=0$, while $A_{2l}^{2l}=a^{(l)}(Z_{t})$. Therefore,
$A_{2l}^{\nabla}A_{2l}=a^{(l)}(Z_{t})\partial_{2l}a^{(l)}(Z_{t})\partial_{z_{2l}}.$
From here we can write $A_{0}$ as
$A_{0}=\begin{pmatrix}b_{1}^{(1)}(Z_{t})\\ b_{2}^{(1)}(Z_{t})\\ ...\\ b_{1}^{(N)}(Z_{t})\\ b_{2}^{(N)}(Z_{t})\end{pmatrix}-\frac{1}{2}\begin{pmatrix}0\\ a^{(1)}(Z_{t})\partial_{x^{1}}a^{(1)}(Z_{t})\\ ...\\ 0\\ a^{(N)}(Z_{t})\partial_{x^{N}}a^{(N)}(Z_{t})\end{pmatrix}=\begin{pmatrix}b_{1}^{(1)}(Z_{t})\\ b_{2}^{(1)}(Z_{t})-\frac{1}{2}a^{(1)}(Z_{t})\partial_{x^{1}}a^{(1)}(Z_{t})\\ ...\\ b_{1}^{(N)}(Z_{t})\\ b_{2}^{(N)}(Z_{t})-\frac{1}{2}a^{(N)}(Z_{t})\partial_{x^{N}}a^{(N)}(Z_{t})\end{pmatrix}$.
We can now compute $A_{0},A_{2k}]$ for k=1,,Nk=1,...,N. By definition of the Lie brackets
@@
Then,
$A_{0}^{\nabla}A_{2k}=\sum_{i=1}^{2N}A_{0}^{i}\partial_{i}a^{(k)}(Z_{t})\partial_{z_{2k}},$
that can be seen as
$\begin{pmatrix}0\\ ...\\ \sum_{i=1}^{2N}A_{0}^{i}\partial_{i}a^{(k)}(Z_{t})\\ ...\\ 0\end{pmatrix}$,
where the only component different from 0 is in position $2k$. Moreover,
$A_{2k}^{\nabla}A_{0}=\sum_{j=1}^{2N}a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{j}\partial_{z_{j}}.$
Let us now introduce, for l=1,,Nl=1,...,N, $f_{l}(Z_{t}):=b_{2}^{(l)}(Z_{t})\frac{1}{2}a^{(l)}(Z_{t})\partial_{x^{l}}a^{(l)}(Z_{t})$. Then,
$A_{0},A_{2k}]=\begin{pmatrix}-a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{1}\\ -a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{2}\\ ...\\ -a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{2k-1}\\ -a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{2k}+\sum_{i=1}^{2N}A_{0}^{i}\partial_{i}a^{(k)}(Z_{t})\\ ...\\ -a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{2N-1}\\ -a^{(k)}(Z_{t})\partial_{x^{k}}A_{0}^{2N}\end{pmatrix}=\begin{pmatrix}-a^{(k)}(Z_{t})\partial_{x^{k}}b_{1}^{(1)}(Z_{t})\\ -a^{(k)}(Z_{t})\partial_{x^{k}}f_{1}(Z_{t})\\ ...\\ -a^{(k)}(Z_{t})\partial_{x^{k}}b_{1}^{(k)}(Z_{t})\\ -a^{(k)}(Z_{t})\partial_{x^{k}}f_{k}(Z_{t})+\sum_{i=1}^{2N}A_{0}^{i}\partial_{i}a^{(k)}(Z_{t})\\ ...\\ -a^{(k)}(Z_{t})\partial_{x^{k}}b_{1}^{(N)}(Z_{t})\\ -a^{(k)}(Z_{t})\partial_{x^{k}}f_{N}(Z_{t})\end{pmatrix}$.
Observe that $A_{0},A_{2k_{1}}]$ and $A_{0},A_{2k_{2}}]$ are independent for $k_{1}\neq k_{2}$, due to the fact that for any $j,k\in\{1,\dots,N\}$ with $j\neq k$, we have $\partial_{x^{k}}b_{1}^{(k)}(z)\neq\partial_{x^{k}}b_{1}^{(j)}(z)\neq 0$.
Consequently, a basis for $\mathbb{R}^{2N}$ is given by $(A_{2k},[A_{0},A_{2k}])_{k\in\{1,\dots,N\}}$. This holds because we have assumed that $\partial_{x^{k}}b_{1}^{(i)}(z)\neq 0$ for all $k,i\in\{1,\dots,N\}$, and that $a^{(k)}(z)$ is non-degenerate for each $k\in\{1,\dots,N\}$.

### v3

@@ Appendix A Proof of technical results
having also used the Lipschitz continuity of $a$ and Lemma 6.1(ii), (iii). A similar bound holds true for the moments of $\alpha_{j+1}^{i}$ which, together with (A.2) and (33), provides
@@
By (A.2) we conclude that $\mathbb{E}[|\mathbb{E}_{j\Delta}[\tilde{\beta}^{i}_{j}\tilde{\xi}^{i}_{j}]|^{k}]\leq C\Delta^{\frac{3}{2}k}$, where $C$ does not depend on $i,j,n,N$. The same argument works for the $k$-th absolute moment of $\mathbb{E}_{j\Delta}[\beta^{i}_{j+1}U^{i}_{j}]$. It remains to deal with
@@
In order to prove (39) for every $\theta$ we use [32, Lemma 9]. We note that $F_{j,k}(\theta)$ is $\mathcal{F}^{N}_{(j+2)\Delta}$-measurable, since so is $\tilde{X}_{(j+1)\Delta}^{i}$. Therefore, we apply [32, Lemma 9] separately to the sum of $F_{j,k}(\theta)$ over even $j$ and that one over odd $j$. However, if we gather all the terms back, then it suffices to prove that, for $k=1,2$,
@@
In order to prove the convergence on the LHS of (49) for every $\theta$, we use [32, Lemma 9] in the same way as in the proof of (39): it suffices to prove that
@@
$\sum_{j=2}^{n-2}F_{j,1}\xrightarrow{d}\mathcal{N}(0,\bar{\Pi}((fa)^{2})),\qquad\sum_{j=1}^{n-2}F_{j,2}\xrightarrow{\mathbb{P}}0,\qquad F_{0}\xrightarrow{\mathbb{P}}0.$ (52)
@@
The second relation in (52) follows by [32, Lemma 9] if we prove that both
@@
$R^{N}_{n,1}\xrightarrow{d}\mathcal{N}(0,\bar{\Pi}(f^{2}a^{4})),\qquad\sum_{k=2}^{5}R^{N}_{n,k}\xrightarrow{\mathbb{P}}0.$ (57)
We first study the convergence of $R^{N}_{n,1}$. We recall that $U^{i}_{j}=\tilde{\xi}^{i}_{j}+\xi^{i}_{j+1}$, where $\tilde{\xi}^{i}_{j}$, $\xi^{i}_{j+1}$ are $\mathcal{F}^{N}_{(j+1)\Delta}$-, $\mathcal{F}^{N}_{(j+2)\Delta}$-measurable, respectively, and independent, and $\mathcal{N}(0,\frac{1}{3})$-distributed. We rewrite
@@
We get that $\mathbb{E}[|\rho_{0}|]\leq C(N\Delta)^{\frac{1}{2}}$, which tends to 0, while the sum of $\rho_{j}$ over $j=1,\dots,n-2$ is the main term. In order to prove that this sum converges in distribution to $\mathcal{N}(0,\bar{\Pi}(f^{2}a^{4}))$, we use a central limit theorem for martingale difference arrays (see, e.g. Theorems 3.2 and 3.4 of [44]). In particular, we need to prove that
@@
follows by [32, Lemma 9] (see also (50)) and by Proposition 6.8. This concludes the proof of (59). Finally, (60) follows from the bound $C\Delta\to 0$ on the $L^{1}$-norm of its LHS (see also (55)).
@@
Furthermore, application of [32, Lemma 9] implies that
@@
by $C\Delta$ and $C\Delta$, respectively. Similarly, application of [32, Lemma 9] implies that $R^{N}_{n,3}\xrightarrow{\mathbb{P}}0$ as $N\Delta\to 0$. Also, it is easy to check that $\mathbb{E}[|R^{N}_{n,4}|]\leq C(N\Delta)^{\frac{1}{2}}$ tends to 0. At last, let us prove that $R^{N}_{n,5}\xrightarrow{\mathbb{P}}0$ as $N\Delta\to 0$. By applying the $L^{1}$ argument again, the proof reduces to that of
@@
as $N\Delta\to 0$, which in turn follows by [32, Lemma 9], because using Proposition 6.4(i) we can bound the $L^{1}$-norm of
@@
In the proof below, we follow the formalism introduced in [70, Section 2.3.2]. Note that since the algebraic vector corresponding to the vector field $A_{2k}$ is given by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{2k}=(0,\dots,a^{(k)}(Z_{t}),\dots,0)^{\top}$, and $a^{(k)}(Z_{t})$ is bounded away from zero, the vectors $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{2k})_{k=1,\dots,N}$ are linearly independent and span an $N$-dimensional subspace of $\mathbb{R}^{2N}$. Our goal is to show that to generate the full tangent space $\mathbb{R}^{2N}$, it is sufficient to include the coordinate vectors of the Lie brackets, denoted by $\widehat{[A_{0},A_{2k}]}$, for $k=1,\dots,N$.
In our framework, the sparsity of the diffusion matrix $\hat{A}$ significantly simplifies the analysis. Recall the definition of the drift vector field:
$A_{0}=B-\frac{1}{2}\sum_{l=1}^{2N}A_{l}^{\nabla}A_{l}.$
Observe that this reduces to:
@@
since $A_{l}^{\nabla}A_{l}=0$ for any odd index $l$ (due to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{2l-1}=\mathbf{0}$). Thus, only the even-indexed terms contribute. Moreover, according to the definition of the covariant derivative in (15), we have:
$A_{2l}^{\nabla}A_{2l}=\sum_{i,j=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}^{i}_{2l}\partial_{z_{i}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{2l}^{j}\partial_{z_{j}},$
where $\partial_{z_{j}}=\mathbf{1}_{\{j=2k-1\}}\partial_{y^{k}}+\mathbf{1}_{\{j=2k\}}\partial_{x^{k}}$. Recall also that for all $i\neq 2l$, we have ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}^{i}_{2l}=0$, while ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{2l}^{2l}=a^{(l)}(Z_{t})$. Therefore, the operator simplifies to:
$A_{2l}^{\nabla}A_{2l}=a^{(l)}(Z_{t})\partial_{x^{l}}a^{(l)}(Z_{t})\partial_{z_{2l}}.$
Let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}$ denote the algebraic column vector containing the coefficients of the vector field $A_{0}$. From the derivation above, we can write ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}$ explicitly as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}=\begin{pmatrix}b_{1}^{(1)}(Z_{t})\\ b_{2}^{(1)}(Z_{t})\\ \vdots\\ b_{1}^{(N)}(Z_{t})\\ b_{2}^{(N)}(Z_{t})\end{pmatrix}-\frac{1}{2}\begin{pmatrix}0\\ a^{(1)}(Z_{t})\partial_{x^{1}}a^{(1)}(Z_{t})\\ \vdots\\ 0\\ a^{(N)}(Z_{t})\partial_{x^{N}}a^{(N)}(Z_{t})\end{pmatrix}=\begin{pmatrix}b_{1}^{(1)}(Z_{t})\\ b_{2}^{(1)}(Z_{t})-\frac{1}{2}a^{(1)}(Z_{t})\partial_{x^{1}}a^{(1)}(Z_{t})\\ \vdots\\ b_{1}^{(N)}(Z_{t})\\ b_{2}^{(N)}(Z_{t})-\frac{1}{2}a^{(N)}(Z_{t})\partial_{x^{N}}a^{(N)}(Z_{t})\end{pmatrix}$.
We can now compute the Lie brackets $A_{0},A_{2k}]$ for $k=1,\dots,N$. By definition:
@@
For the first term, we have:
$A_{0}^{\nabla}A_{2k}=\sum_{i=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{i}\partial_{z_{i}}a^{(k)}(Z_{t})\partial_{z_{2k}},$
which corresponds to the algebraic vector:
$\widehat{A_{0}^{\nabla}A_{2k}}=\begin{pmatrix}0\\ \vdots\\ \sum_{i=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{i}\partial_{z_{i}}a^{(k)}(Z_{t})\\ \vdots\\ 0\end{pmatrix}$,
where the only non-zero component is in the $2k$-th position. For the second term, we have:
$A_{2k}^{\nabla}A_{0}=\sum_{j=1}^{2N}a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{j}\partial_{z_{j}}.$
Let us introduce, for $l=1,\dots,N$, the function $f_{l}(Z_{t}):=b_{2}^{(l)}(Z_{t})-\frac{1}{2}a^{(l)}(Z_{t})\partial_{x^{l}}a^{(l)}(Z_{t})$. The coordinate vector for the Lie bracket, $\widehat{[A_{0},A_{2k}]}$, is then given by:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widehat{[A_{0},A_{2k}]}}=\begin{pmatrix}-a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{1}\\ -a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{2}\\ \vdots\\ -a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{2k-1}\\ -a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{2k}+\sum_{i=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{i}\partial_{z_{i}}a^{(k)}(Z_{t})\\ \vdots\\ -a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{2N-1}\\ -a^{(k)}(Z_{t})\partial_{x^{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{2N}\end{pmatrix}=\begin{pmatrix}-a^{(k)}(Z_{t})\partial_{x^{k}}b_{1}^{(1)}(Z_{t})\\ -a^{(k)}(Z_{t})\partial_{x^{k}}f_{1}(Z_{t})\\ \vdots\\ -a^{(k)}(Z_{t})\partial_{x^{k}}b_{1}^{(k)}(Z_{t})\\ -a^{(k)}(Z_{t})\partial_{x^{k}}f_{k}(Z_{t})+\sum_{i=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{0}^{i}\partial_{z_{i}}a^{(k)}(Z_{t})\\ \vdots\\ -a^{(k)}(Z_{t})\partial_{x^{k}}b_{1}^{(N)}(Z_{t})\\ -a^{(k)}(Z_{t})\partial_{x^{k}}f_{N}(Z_{t})\end{pmatrix}$.
Observe that the vectors $^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widehat{[A_{0},A_{2k_{1}}]}}$ and $^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widehat{[A_{0},A_{2k_{2}}]}}$ are linearly independent for $k_{1}\neq k_{2}$. This follows from the assumption that for any $j,k\in\{1,\dots,N\}$ with $j\neq k$, we have $\partial_{x^{k}}b_{1}^{(k)}(z)\neq\partial_{x^{k}}b_{1}^{(j)}(z)\neq 0$.
Consequently, the basis for $\mathbb{R}^{2N}$ is given by the algebraic vectors $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\hat{A}_{2k},\widehat{[A_{0},A_{2k}]})_{k\in\{1,\dots,N\}}$}. This holds precisely because we assumed $\partial_{x^{k}}b_{1}^{(i)}(z)\neq 0$ for all $k,i\in\{1,\dots,N\}$, and that the diffusion coefficient $a^{(k)}(z)$ is non-degenerate for each $k\in\{1,\dots,N\}$. Thus, Hörmander’s condition is satisfied, and the proof is complete.